OpenStudy (landon34):

Tracy is saving to buy a new guitar that costs $253.77 (including tax). Tracy has saved $136.27, and she can save $23.50 each week. Which equation could you use to find how long it will take to save for the guitar? $23.50w = $253.77 $253.77 - w = $136.27 + $23.50 $253.77 + $136.27 = $23.50w $136.27 + $23.50w = $253.77

OpenStudy (md152727):

She will earn $23.5*w in w weeks. We form equation Total Cost = Current savings + future savings 253.77 = 136.27 + 23.5w 253.77-136.27 = 136.27 + 23.5w - 136.27 117.5 = 23.5w 117.5/23.5 = 23.5w/23.5 5 = w w = 5 weeks

OpenStudy (landon34):

ok, thank you.

OpenStudy (md152727):

You welcome